OpenStudy (anonymous):

For the function f(x)=1/x, what is the equation of the tangent line at x=1.

OpenStudy (anonymous):

f ' (x) = -1/x^2 f ' (1) = -1 = Slope f (1) = 1 giving us coordinate (1,1) So our slope is Y-1=-1(X-1)

OpenStudy (anonymous):

simplified makes it y=-x+2?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

thanks!